1.) 75/0.006
2.) square root of 0.0016
On these i need to change them to decimals and i can't use a caculator, need some help, thank you in advance for help.
1.) 1/12
2.) 3/8

Answer by stanbon(48553)

1.) 75/0.006
= 75/(6/1000)
= 25/(2/1000)
= 25/(1/500)
= 25*500
= 12500
=================
2.) square root of 0.0016
sqrt(16/10000)
= 4/100
= 1/25 or 0.04
-------------------
On these i need to change them to decimals and i can't use a caculator, need some help, thank you in advance for help.
1.) 1/12
Divide 12 into 1.00000 to get 0.0833
2.) 3/8
Divide 8 into 3.00000 to get 0.375
================
Cheers,
Stan H.